Night Plan UT 01-Dec-2020

  • Starting Configuration LUCI-Bino in the beam, LUCI-1 authorized
  • AO Science and Technical
    • DONE OSCO Checkout LUCI-1 AO at Twilight - most any conditions
    • Follwing Require Clear (or very thin clouds) and good seeing:
      • DONE GRB 200826A - AO H-band observations (2hrs) - Rothberg (rollover from last D nights) - clear preferred, thin cirrus OK, DIMM 1" or better - 1st half of night
        • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/GRB_200826A
        • Observe PSF star along with science target - AORef star manually set to R=16.51 to force Bin 4
      • UDS z~10 target - ESM/L1 spectroscopy w/L2 (SL) in parallel (3 hrs) - Rothberg (rollover from last D nights) - clear preferred, thin cirrus acceptable DIMM 1" or better - 1st half of night
        • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/UDS_z10_ESM_Science
        • Observations are mixed-mode: L1-ESM and L2-SL. Same setup on both sides (N3.75, G210 HKspec, CWL=1.6, 0.75" slit)
        • Blind offset along slit - center on AO Ref first then offset to faint galaxy - target is H=25.3 (won't see in acq images). AO Ref star remains in slit during observations
        • AO Ref R=14
    • If UDS z~10 cannot be executed, run the Mars AO script : Mars_AO_UTDec01_UT0400_L1.xml
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Mars_AO
      • Mars will be 40 degrees from Moon
      • Can be executed in thin clouds, DIMM up to 1.2"
      • Must be started by UT 04:00 or 9pm local time
      • Scripts are hand editd to allow H2+ND2 combination
      • Non-Sidereal Target: Bring up NSIGUI, search for Phobos (this is our AO Ref Star)
      • Insert Pause after first exposure - this is the acquisition step - it uses FeII +ND2
      • Use the TELSVC to center Mars in the LUCI-1 N30 FOV
      • Be sure to use absorb offsets in TELSVC after Mars is centered up
      • Continue with Script
    • If UDS or Mars cannot be executed and its clear - go to Palomar 2 (new version) -
      • If photometric/clear - take standard before and after (P247U - this cycles through all 5 broad band filters)
      • If DIMM < 1.2" but clouds - still execute
      • New Version of script has PA=0 and different dithers to avoid hitting baystage limits - same AO Ref star (double but separable with OCAM)
    • UT_dec01_targets.jpg
  • DONE Midnight re-configure for LBCs (rollover of previous night's work)
    • Filter focus offset tests - 2hrs - clear/thin skies - requires seeing 1" or better
    • TMS testing - 2-3hrs
    • non-lineariry/gain tests - 1 hr - requires photometric conditions
  • Poor Weather/Poor Seeing Condition Programs
    • ESM data collection - if you can close the loop take the data
      • M52 (3 filtersand bright and faint AO Ref star) - highest priority for first half of night - visible at twilight hits limit at 1am local time
      • M34 (3 filters) - medium priority - transits at 10pm - visible at twlight evening
      • M67 (3 filters and bright and faint AO Ref star)
    • ICRF3 Spectroscopy (Rothberg)
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Poor_Weather/ICRF3_Rothberg_2020B
      • Observe telluric first
      • Double Check high Dec objects to make sure tellurc/science are close spatially
    • If clear but poor seeing, please observe standard stars, cycling through z,J,H,K,Ks
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Poor_Weather/LUCI_Phot_ZP
      • Select stars to use based on visibility - OT file is in directory, export the scripts as needed
      • Please observe the same star at least two times at different airmasses

Nightlog link

-- BarryRothberg - 30 Nov 2020
Topic revision: r4 - 15 Dec 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback